(1 point) Represent the function 82 2+2 as as a power series f(z) = { Co 8x Il C2 cs CA Find the radius of convergence R =
Answer #1

The given function is f(u) = 84 2tu =8n 2(1+4) -1 = 41(1+u 244 [-)+(-()+(*). --------| 1+t) - 1- t + t² - ² + tl. น + = 44- 2l cu Cuti R=lim ol eight! 4.2 (1)ht? 4 zhim ann lim na 1-21=2 ::R=2